OpenStudy (anonymous):

how to understand that an inductor behaves like short circuit and capacitor like an open circuit in steady state in a network

OpenStudy (anonymous):

It depends on the steady-state frequency. At DC (zero frequency) an inductor behaves like a short circuit and a capacitor behaves like an open circuit. As the frequency increases, the inductor acts more and more like an open circuit and a capacitor acts more and more like a short circuit. The physics is pretty easy to see for a capacitor.Think of the capacitor as two metal plates separated by a dielectric. Apply a voltage across the capacitor. The plates build up a charge Q = CV, where C is the capacitance (in farads) and V is the applied voltage. (The capacitor is storing energy in an electric field.) Differentiating both sides with respect to time (and assuming the capacitance remains constant) we get: dQ/dt = CdV/dt But dQ/dt is just the change in charge with respect to time, which is defined as the current (I); so, dQ/dt = I = CdV/dt. So, suppose we connect a sine wave voltage source across the capacitor: \[V=V _{0}\sin \omega t\] The current is then: \[I=\omega C V _{0}\cos \omega t\] Look at the ratio of the amplitude of the voltage divided by the amplitude of the current: \[V/I=1/\omega C\] This is the "resistance" (actually called "impedance") of the capacitor. When \[\omega = 0\] the impedance becomes infinite (an open circuit). On the other hand, if the frequency gets very high, the impedance gets very low. The inductor stores energy in a magnetic field, which is proportional to the current through the inductor times the value of inductance. For an inductor with an inductance of L henrys, the V-I relation is V = L dI/dt. Suppose we have a circuit with a current source driving an inductor with inductance L, where the source current is given by: \[I=I _{0}\sin \omega t\] Then the voltage across the inductor is \[V=\omega LI _{0}\cos \omega t\] Now the ratio of the amplitudes, i.e., the effective impedance, is given by: \[V/I=\omega L\] So at zero frequency (DC), the inductance has zero impedance (a short circuit) Note that in the second example it was specified that the inductance was driven by a current source to allow for the possibility of direct current. A good way to blow out a transformer, which is effectively two coupled inductors, would be to connect a DC voltage source across one of the windings. So to recap, the answer to your question is that the steady-state behavior of capacitors and inductors depends on the steady-state frequency: - at DC (zero frequency), an inductor acts like a short circuit and a capacitor acts like an open circuit - at infinite frequency, an inductor acts like an open circuit and a capacitor acts like a short circuit

OpenStudy (anonymous):

thanks a lot